May I have some help with this inequality?
Solve the nonlinear inequality. Express the solution using interval notation. (If the solution contains only one interval, use (0, 0) for the remaining intervals. If you need to enter - or , type -INFINITY or INFINITY.)
x^2 + 4x ≤ 60

Answer by checkley71(8403)

x^2+4x<=60
x^2+4x-60<=0
(x+10)(x-6)<=0
x+10=0
x=-10 answer.
x-6=0
x=6 answer.